# These authors contributed equally

### v2

@@ 1 Introduction
In the era of large and complex neural models, ensuring trust between them and humans becomes a critical issue. Explainability literature suggests a variety of methods to probe neural models behaviors, either requiring access to their inner workings [11, 29] or not [6, 7, 10]. There is also ample discussion regarding the nature of the features involved in a human-understandable explanation; for example, low-level features e.g. pixel-related characteristics (brightness, contrast) may be unable to provide a meaningful explanation to the end user, despite being informative for a neural model [27]. This observation is applicable in the counterfactual explanation scenario, suggesting that semantics are fundamentally essential to meaningful counterfactual explanations [4]. In the meanwhile, both high-level semantics as well as low-level features can be expressed in the same mathematical format (e.g. numerical vectors), allowing the transition to semantically rich explanation systems [6]. In the interest of explaining visual classification systems, and with respect to recent related work [2, 6, 7, 10, 11, 29], we base our current study on conceptual counterfactual explanations.
@@
With regard to the model accessibility an explainability system can have, most related counterfactual explanation methods belong to the white-box category [2, 11, 29], offering insights tailored to the model under investigation. However, in the emergence of powerful, yet proprietary models, such as ChatGPT [23] and GPT-4 [24] which are currently accessible only through APIs, black-box explanations resurface as more viable and universal solutions, able to approach a model by scrutinizing their outputs. Under these constrains, we stay within the black-box setting of conceptual counterfactuals [6, 7, 10].
We also acknowledge the importance of intra-concept roles, as associations between the same concepts via roles may result in different interpretations of situations (Figure 1). Thus, we advocate image representation via scene graphs. To this end, we extend the framework of [7] to accommodate a larger spectrum of Graph Machine Learning algorithms to serve conceptual counterfactual explanations on scene images. We retain the brute-force calculation of Graph Edit Distance (GED)[28] as the ground truth counterfactual measure, under which the closest scene graph in terms of GED belonging to a different class is considered as the ”gold” counterfactual graph. In order to overcome the computational complexity of GED calculation, we employ Graph Machine Learning algorithms, such as Graph Kernels, Graph Autoencoders (GAEs) and supervised Graph Neural Networks (GNNs), which are proven to be lightweight and significantly faster than the deterministic GED calculation, even when optimizations [14] are employed. Graph edits serving as counterfactual explanations are instructed via concept distances in existing hierarchies, in our case being WordNet [21]. The outline of our approach is presented in Figure 2. Overall, we explore the capabilities of Graph Machine Learning algorithms, contributing to the following:
@@ 3 Method
The focus of our current study centers on determining the minimal semantic alteration necessary for an image $i$ to transition from being classified as A to B. Each image $i$ in our dataset $I$ is represented by a scene graph $G_{i}$, encapsulating semantic information, therefore forming an explanation dataset {$i,G_{i}$}, as previously outlined in [10]. Utilizing a reference image $i_{r}$ and candidate scene graphs $G_{i_{c}}$, the Graph Edit Distance (GED) is computed using an optimization approach to identify the most efficient path for semantic transformation between $G_{i_{r}}$ and $G_{i_{c}}$, guided by predefined graph edit operations (Replacement, Deletion, Insertion). Semantic distances, drawn from the well-crafted WordNet [21] hierarchy, guide the choice of edits, ensuring meaningful transitions while penalizing irrelevant replacements. To accelerate GED calculations, we employ optimization techniques similar to [7], utilizing the Volgenant-Jonker (VJ) algorithm [14]. Finally, to obtain the counterfactual candidate $G_{i_{cc}}$ of the original scene graph $G_{i_{r}}$, we repeat the optimized GED calculation for each candidate $G_{i_{c}}$, where $c\in C=I\setminus\{i_{r}\}$, ensuring that the selected closest $G_{i_{cc}}$ belongs to class $B\neq A$, while $G_{i_{r}}$ belongs to class $A$, formulated as:
@@ 4.1 Experimental setup
In our study, we employ various graph kernels [22] including Shortest Path (SP), Weisfeiler-Lehman (WL), Neighborhood Hash (NH), Random Walk (RW), and Graphlet Sampling (GS), implemented using the GraKel library. For supervised Graph Neural Network (GNN) models, we utilize single-layer siamese Graph Convolutional Network (GCN), Graph Attention Network (GAT), and Graph Isomorphism Network (GIN) as described in [7], trained over 50 epochs with a batch size of 32 and Adam optimizer. Regarding Graph Autoencoders (GAEs), we implement vanilla GAE, Variational GAE (VGAE) [16], and Adversarially Regularized Variational Graph Autoencoder (ARVGA) [25], alongside a variation of Graph Feature Autoencoder (GFA) [12]. GAE and VGAE are both based on the generic auto-encoder structure, but differ in terms of loss function. ARVGA employs basic concepts from Generative Adversarial Networks to further boost the regularization of the latent embeddings produced by VGAE. Finally, GFA utilizes the Feature Decoder (whose goal is to predict the feature matrix $X$), alongside the original VGAE Inner-Product Decoder (whose goal is to predict the adjacency matrix $A$). All these models incorporate single-layer GCN, GAT, and GIN architectures in the encoder (and the decoder for GFA), employing AdamW optimizer with default parameters and a learning rate of 0.001, trained for 20 epochs with a batch size of 32.
GNNs and GAEs were implemented using PyTorch Geometric [9]. In all cases, cosine similarity was used as a measure for embedding similarity. For each scene graph in the dataset acting as the query, we retrieve the top $k$ closest graphs with different PLACES-365 labels using a pretrained PLACES classifier with a ResNet50 backbone, resulting in slightly improved labels compared to [7]. We then evaluate whether the counterfactual graph, determined by the ground truth Graph Edit Distance (GED), is found within this ranked list of $k$ items.
@@
Information retrieval metrics are used for evaluation, specifically Precision (P@k) and Normalized Discounted Cumulative Gain (NDCG@k). P@k returns the percentage of the relevant items, when considering the first $k$ items in the rank. We also employ the ”binary” variant from [7], where only the top-1 Graph Edit Distance (GED) item is relevant, in line with the counterfactual logic. Binary NDCG@k compares the rank of the top-1 counterfactual returned by an algorithm to the ideal ground truth rank obtained by GED. While NDCG@k lacks direct interpretability, it’s widely used for comparing retrieval system performance. Our evaluation focuses on $k$=1, 2, 4 ranks, once again emphasizing proximity to the top-1 result. Additional metrics assess the proposed edits’ minimality by GNNs, including average node, edge, and overall edits, alongside average GED of counterfactuals. Lower values indicate better model performance, with these metrics considered relative to each other.
@@ 4.3 Qualitative Results
In the case of VG-DENSE, multiple instances exhibit concordance among models, owing to its dense interconnections and the inherent ability of all models to recognize matching structures. The 1st row demonstrates unanimous agreement, as the obtained scene graph essentially forms a subgraph of the query. A human could also agree on the semantic closeness between counterfactual instances: Semantics such as ”table”, ”cup”, ”phone” are preserved, while the kitchen-related ”cake” and ”newspaper” should be edited to allow workstation-related semantics, such as ”laptop”. Moving to the 2nd row, GNN methods align; successfully integrating structural and semantic information. Despite the kernel-retrieved result having a more similar structure, GNN methods successfully discern and retrieve an image featuring a cat on furniture, surpassing a mere depiction of a bed. In terms of human perception, counterfactuals depicting a ”bedroom” (query) and a ”living room” instance (GNN-based retrievals) respect the change of class label (”bedroom”$\rightarrow$”living room”) with minimally altering semantics (”pillow” and ”cat” semantics are preserved), allowing effortless high-level interpretability. Conversely, in the 3rd row with a widespread disagreement, the merits of the supervised GNN network become evident. The unsupervised GAE captures an image resembling the query with a teddy bear wearing a bow but overlooks the presence of a dog. In contrast, the supervised GNN, having more accurately approximated GED, retrieves an image centered on a dog and a different type of toy. Interestingly, the kernel fails and retrieves an image of a train, emphasizing that a sole focus on structure (star subgraph) is insufficient.
Assessing retrieved VG-RANDOM results proves challenging due to the dissimilar nature of the sparser underlying graphs. In the 4th row, the supervised GCN retrieves a structurally closer scene graph, despite visual similarities in GNN-based counterfactuals. It becomes harder even for humans to evaluate their proximity, since all instances contain ”chair” and ”curtain” concepts, but no other discriminative structure or detail. In the final row the models reach consensus, successfully retrieving an image with palm trees, a characteristic the kernel fails to achieve despite the retrieval of very closely aligned graphs. The effectiveness of the framework relies on rich interconnections observed in VG-DENSE, underscoring the significance of expressive scene graphs, human intuition, and thoughtful data preprocessing for enhancing explanation informativeness.
@@ 5 Conclusion
In this work, we unify graph-based conceptual counterfactual explanation algorithms under a common framework, upon which we conduct a comparative study. Various graph-based algorithms, such as kernels, supervised and unsupervised GNNs, highlight the value of graph representations for counterfactuals, offering insights on what needs to be changed conceptually to transit to another class efficiently. Quantitative and qualitative analysis showcases the different criteria that influence the retrieval of counterfactual instances from varying graph algorithms, and the importance of well-defined and informative semantics towards human-interpretable explanations.

### v3

@@ 1 Introduction
In the era of large and complex neural models, ensuring trust between them and humans becomes a critical issue. Explainability literature suggests a variety of methods to probe neural models behaviors, either requiring access to their inner workings [11, 31] or not [6, 7, 10]. There is also ample discussion regarding the nature of the features involved in a human-understandable explanation; for example, low-level features e.g. pixel-related characteristics (brightness, contrast) may be unable to provide a meaningful explanation to the end user, despite being informative for a neural model [29]. This observation is applicable in the counterfactual explanation scenario, suggesting that semantics are fundamentally essential to meaningful counterfactual explanations [4, 23]. In the meanwhile, both high-level semantics as well as low-level features can be expressed in the same mathematical format (e.g. numerical vectors), allowing the transition to semantically rich explanation systems [6]. In the interest of explaining visual classification systems, and with respect to recent related work [2, 6, 7, 10, 11, 31], we base our current study on conceptual counterfactual explanations.
@@
With regard to the model accessibility an explainability system can have, most related counterfactual explanation methods belong to the white-box category [2, 11, 31], offering insights tailored to the model under investigation. However, in the emergence of powerful, yet proprietary models, such as ChatGPT [25] and GPT-4 [26] which are currently accessible only through APIs, black-box explanations resurface as more viable and universal solutions, able to approach a model by scrutinizing their outputs. Under these constraints, we stay within the black-box setting of conceptual counterfactuals [6, 7, 10].
We also acknowledge the importance of intra-concept roles, as associations between the same concepts via roles may result in different interpretations of situations (Figure 1). Thus, we advocate image representation via scene graphs. To this end, we extend the framework of [7] to accommodate a larger spectrum of Graph Machine Learning algorithms to serve conceptual counterfactual explanations on scene images. We retain the brute-force calculation of Graph Edit Distance (GED)[30] as the ground truth counterfactual measure, under which the closest scene graph in terms of GED belonging to a different class is considered as the “gold” counterfactual graph. In order to overcome the computational complexity of GED calculation, we employ Graph Machine Learning algorithms, such as Graph Kernels, Graph Autoencoders (GAEs) and supervised Graph Neural Networks (GNNs), which are lightweight and significantly faster than the NP-hard deterministic GED calculation [35], even when optimizations are employed [14]. Graph edits serving as counterfactual explanations are instructed via concept distances in existing hierarchies, in our case being WordNet [22]. Merits of such explanations include but are not limited to actionability, a crucial quality of counterfactuals in several domains [21]. The outline of our approach is presented in Figure 2. Overall, we explore the capabilities of Graph Machine Learning algorithms, contributing to the following:
@@ 3 Method
The focus of our current study centers on determining the minimal semantic alteration necessary for an image $i$ to transition from being classified as A to B. Each image $i$ in our dataset $I$ is represented by a scene graph $G_{i}$, encapsulating semantic information, therefore forming an explanation dataset {$i,G_{i}$}, as previously outlined in [10]. Utilizing a reference image $i_{r}$ and candidate scene graphs $G_{i_{c}}$, the GED is computed using an optimization approach to identify the most efficient path for semantic transformation between $G_{i_{r}}$ and $G_{i_{c}}$, guided by predefined graph edit operations (Replacement, Deletion, Insertion). Semantic distances, drawn from the well-crafted WordNet [22] hierarchy, guide the choice of edits, ensuring meaningful transitions while penalizing irrelevant replacements. To accelerate GED calculations, we employ optimization techniques similar to [7], utilizing the Volgenant-Jonker (VJ) algorithm [14]. Finally, to obtain the counterfactual candidate $G_{i_{cc}}$ of the original scene graph $G_{i_{r}}$, we repeat the optimized GED calculation for each candidate $G_{i_{c}}$, where $c\in C=I\setminus\{i_{r}\}$, ensuring that the selected closest $G_{i_{cc}}$ belongs to class $B\neq A$, while $G_{i_{r}}$ belongs to class $A$, formulated as:
@@ 4.1 Experimental setup
In our study, we employ various graph kernels [24] including Shortest Path (SP), Weisfeiler-Lehman (WL), Neighborhood Hash (NH), Random Walk (RW), and Graphlet Sampling (GS), implemented using the GraKel library. For supervised Graph Neural Network (GNN) models, we utilize single-layer siamese GCN, GAT, and GIN as described in [7], trained over 50 epochs with a batch size of 32 and Adam optimizer. Regarding Graph Autoencoders (GAEs), we implement vanilla GAE, Variational GAE (VGAE) [16], and Adversarially Regularized Variational Graph Autoencoder (ARVGA) [27], alongside a variation of Graph Feature Autoencoder (GFA) [12]. GAE and VGAE are both based on the generic auto-encoder structure, but differ in terms of loss function. ARVGA employs basic concepts from Generative Adversarial Networks to further boost the regularization of the latent embeddings produced by VGAE. Finally, GFA utilizes the Feature Decoder (whose goal is to predict the feature matrix $X$), alongside the original VGAE Inner-Product Decoder (whose goal is to predict the adjacency matrix $A$). All these models incorporate single-layer GCN, GAT, and GIN architectures in the encoder (and the decoder for GFA), employing AdamW optimizer with default parameters and a learning rate of 0.001, trained for 20 epochs with a batch size of 32.
GNNs and GAEs were implemented using PyTorch Geometric [9]. In all cases, cosine similarity was used as a measure for embedding similarity. For each scene graph in the dataset acting as the query, we retrieve the top $k$ closest graphs with different PLACES-365 labels using a pretrained PLACES classifier with a ResNet50 backbone, resulting in slightly improved labels compared to [7]. We then evaluate whether the counterfactual graph, determined by the ground truth GED, is found within this ranked list of $k$ items.
@@
Information retrieval metrics are used for evaluation, specifically Precision (P@k) and Normalized Discounted Cumulative Gain (NDCG@k). P@k returns the percentage of the relevant items, when considering the first $k$ items in the rank. We also employ the “binary” variant from [7], where only the top-1 GED item is relevant, in line with the counterfactual logic. Binary NDCG@k compares the rank of the top-1 counterfactual returned by an algorithm to the ideal ground truth rank obtained by GED. While NDCG@k lacks direct interpretability, it’s widely used for comparing retrieval system performance. Our evaluation focuses on $k$=1, 2, 4 ranks, once again emphasizing proximity to the top-1 result. Additional metrics assess the proposed edits’ minimality by GNNs, including average node, edge, and overall edits, alongside average GED of counterfactuals. Lower values indicate better model performance, with these metrics considered relative to each other.
@@ 4.3 Qualitative Results
In the case of VG-DENSE, multiple instances exhibit concordance among models, owing to its dense interconnections and the inherent ability of all models to recognize matching structures. The 1st row demonstrates unanimous agreement, as the obtained scene graph essentially forms a subgraph of the query. A human could also agree on the semantic closeness between counterfactual instances: Semantics such as “table”, “cup”, “phone” are preserved, while the kitchen-related “cake” and “newspaper” should be edited to allow workstation-related semantics, such as “laptop”. Moving to the 2nd row, GNN methods align; successfully integrating structural and semantic information. Despite the kernel-retrieved result having a more similar structure, GNN methods successfully discern and retrieve an image featuring a cat on furniture, surpassing a mere depiction of a bed. In terms of human perception, counterfactuals depicting a “bedroom” (query) and a “living room” instance (GNN-based retrievals) respect the change of class label (“bedroom”$\rightarrow$“living room”) with minimally altering semantics (“pillow” and “cat” semantics are preserved), allowing effortless high-level interpretability. Conversely, in the 3rd row with a widespread disagreement, the merits of the supervised GNN network become evident. The unsupervised GAE captures an image resembling the query with a teddy bear wearing a bow but overlooks the presence of a dog. In contrast, the supervised GNN, having more accurately approximated GED, retrieves an image centered on a dog and a different type of toy. Interestingly, the kernel fails and retrieves an image of a train, emphasizing that a sole focus on structure (star subgraph) is insufficient.
Assessing retrieved VG-RANDOM results proves challenging due to the dissimilar nature of the sparser underlying graphs. In the 4th row, the supervised GCN retrieves a structurally closer scene graph, despite visual similarities in GNN-based counterfactuals. It becomes harder even for humans to evaluate their proximity, since all instances contain “chair” and “curtain” concepts, but no other discriminative structure or detail. In the final row the models reach consensus, successfully retrieving an image with palm trees, a characteristic the kernel fails to achieve despite the retrieval of very closely aligned graphs. The effectiveness of the framework relies on rich interconnections observed in VG-DENSE, underscoring the significance of expressive scene graphs, human intuition, and thoughtful data preprocessing for enhancing explanation informativeness.
@@ 5 Conclusion
In this work, we unify graph-based conceptual counterfactual explanation algorithms under a common framework, upon which we conduct a comparative study. Various graph-based algorithms, such as kernels, supervised and unsupervised GNNs, highlight the value of graph representations for counterfactuals, offering insights on what needs to be changed conceptually to transit to another class efficiently. Quantitative and qualitative analysis showcases the different criteria that influence the retrieval of counterfactual instances from varying graph algorithms, and the importance of well-defined and informative semantics towards human-interpretable explanations. There is ample room for further real-world applications of this framework within various domains where we have access to graph-structured data, such as recommendation systems, transportation networks, as well as agricultural and health inspection sectors.
{credits}
5.0.1 Acknowledgements
This research work is Co-funded from the European Union’s Horizon Europe Research and Innovation programme under Grant Agreement No 101119714 — dAIry 4.0.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.